# A simple experiment for observing clustering and dynamics of coalescing particles in air turbulence

### 09 Mar 2026 v1

@@ 2.3 Particle size characterization
To obtain accurate particle diameters, a physically grounded intensity-thresholding scheme was applied. In this kind of imaging setup, which some would call ”shadow imaging”, droplets appear darker than the background in the raw images (12-bits/pixel), a lower intensity corresponds to a more focused particle. The mean background pixel intensity $\langle I_{\rm b}\rangle$ was estimated by analyzing the pixel intensities of particle free regions in each frame. In this section, $\langle\cdot\rangle$ denotes a spatial average over particle free pixels within a single frame. These regions were identified in a two-step procedure. First, a preliminary standard deviation $\sigma_{\mathrm{0}}$ was estimated directly from the raw intensity field of the image. At this stage, the statistics include both background illumination fluctuations and particle-induced intensity deficits. Pixels with intensity below $3\sigma_{\mathrm{0}}$ were excluded to remove particle-containing pixels. Based on the remaining particle-free pixels, the background intensity fluctuations (standard deviation) $\sigma_{b}$ were subsequently evaluated. The resulting $\sigma_{b}$, was estimated to be $0.014$, $0.013$, and $0.012$ for experiment with spinning-disk speed of 42,000 rpm, 30,000 rpm, and 18,000 rpm, respectively. Here, the image intensity $I$ was normalized to the range $[0,1]$. Incorporating these noise estimates, a particle detection threshold $T_{\rm 1}$ was defined as
@@ 2.4 Turbulence characterization and particle dynamic parameters
The turbulent flow in the measurement region is characterized by its velocity fluctuations, Reynolds number, energy dissipation rate, and the associated Kolmogorov scales. The turbulence intensity is quantified by the root-mean-square velocity fluctuation $u_{\mathrm{rms}}$, from which the Reynolds number $Re$ is evaluated.
@@
where $\nu$ is the kinematic viscosity of air, and $L$ is the characteristic length scale of the flow. In the present experiments, $L=0.1~\mathrm{m}$ is adopted based on empirical considerations (i.e., it is taken as half of the chamber width). Here, $Re$ is introduced as a nominal Reynolds number and is used solely for comparing different operating conditions rather than for formal theoretical analysis.
@@
Once $\tau_{\eta}$ is determined, the Stokes number ($St$) is defined as the ratio of the particle dynamic response time ($\tau_{p}$) to the Kolmogorov time scale ($\tau_{\eta}$):
@@
In our experiments, the ambient temperature varied between 18 °C and 27 °C. The temperature dependence of the fluid density is negligible within this range: the variation in air density is less than 3%, and the densities of water and solid particles remain essentially constant. Therefore, only the temperature-induced change in air kinematic viscosity is corrected. Taking $\nu_{20}=1.48\times 10^{-5}\,\mathrm{m^{2}/s}$ as the reference value at 20 °C, the corresponding viscosity variation is $\nu=(1.51\pm 0.06)\times 10^{-5}\,\mathrm{m^{2}/s}$, which is accounted in the evaluation of uncertainty in $\tau_{p}$ and $St$.
@@
The particle properties and flow parameters for all experiments are summarized in Table 3, including the particle diameter $d$, turbulent velocity fluctuation $u_{\mathrm{rms}}$, Reynolds number $Re$, and particle response time $\tau_{p}$. Based on these measurements, turbulence parameters inferred from both the RDF-based method and the second-order velocity structure function ($S_{2}$) method are listed, including the dissipation rate $\varepsilon$, Kolmogorov length scale $\eta$, and the corresponding Stokes number $St$.
@@ 2.5 Computing the Radial Distribution Function
where $\psi(r)$ represents the summed number of particles located at a separation $r$ from each trial particle (the trial particles are taken to be the entire population of $N$ particles, double counting is allowed). The quantity $\delta V_{r}/V$ denotes the ratio of a small sampling volume at distance $r$ (i.e., a spherical shell of thickness $\delta r$) to the total measurement volume $V$.

### 11 Mar 2026 v2

@@ 2.3 Particle size characterization
To obtain accurate particle diameters, a physically grounded intensity-thresholding scheme was applied. In this kind of imaging setup, which some would call ”shadow imaging”, droplets appear darker than the background in the raw images (12-bits/pixel), a lower intensity corresponds to a more focused particle. The mean background pixel intensity $\langle I_{\rm b}\rangle$ was estimated by analyzing the pixel intensities of particle free regions in each frame. In this section, $\langle\cdot\rangle$ denotes a spatial average over particle free pixels within a single frame. These regions were identified in a two-step procedure. First, a preliminary standard deviation $\sigma_{\mathrm{0}}$ was estimated directly from the raw intensity field of the image. At this stage, the statistics include both background illumination fluctuations and particle-induced intensity deficits. Pixels with intensity below $3\sigma_{\mathrm{0}}$ were excluded to remove particle-containing pixels. Based on the remaining particle-free pixels, the background intensity fluctuations (standard deviation) $\sigma_{b}$ were subsequently evaluated. The resulting $\sigma_{b}$, was estimated to be $0.014$, $0.013$, and $0.012$ for experiment with spinning-disk speed of 42,000 rpm, 30,000 rpm, and 18,000 rpm, respectively. Here, the image intensity II was normalized to the range $[0,1]$. Incorporating these noise estimates, a particle detection threshold $T_{\rm 1}$ was defined as
@@ 2.4 Turbulence characterization and particle dynamic parameters
The turbulent flow in the measurement region is characterized by its velocity fluctuations, Reynolds number, energy dissipation rate, and the associated Kolmogorov scales. The turbulence intensity is quantified by the root-mean-square velocity fluctuation $u_{\mathrm{rms}}$, from which the Reynolds number ReRe is evaluated.
@@
where $\nu$ is the kinematic viscosity of air, and $L$ is the characteristic length scale of the flow. In the present experiments, $L=0.1~\mathrm{m}$ is adopted based on empirical considerations (i.e., it is taken as half of the chamber width). Here, ReRe is introduced as a nominal Reynolds number and is used solely for comparing different operating conditions rather than for formal theoretical analysis.
@@
Once $\tau_{\eta}$ is determined, the Stokes number (StSt) is defined as the ratio of the particle dynamic response time ($\tau_{p}$) to the Kolmogorov time scale ($\tau_{\eta}$):
@@
In our experiments, the ambient temperature varied between 18 °C and 27 °C. The temperature dependence of the fluid density is negligible within this range: the variation in air density is less than 3%, and the densities of water and solid particles remain essentially constant. Therefore, only the temperature-induced change in air kinematic viscosity is corrected. Taking $\nu_{20}=1.48\times 10^{-5}\,\mathrm{m^{2}/s}$ as the reference value at 20 °C, the corresponding viscosity variation is $\nu=(1.51\pm 0.06)\times 10^{-5}\,\mathrm{m^{2}/s}$, which is accounted in the evaluation of uncertainty in $\tau_{p}$ and StSt.
@@
The particle properties and flow parameters for all experiments are summarized in Table 3, including the particle diameter $d$, turbulent velocity fluctuation $u_{\mathrm{rms}}$, Reynolds number ReRe, and particle response time $\tau_{p}$. Based on these measurements, turbulence parameters inferred from both the RDF-based method and the second-order velocity structure function ($S_{2}$) method are listed, including the dissipation rate $\varepsilon$, Kolmogorov length scale $\eta$, and the corresponding Stokes number StSt.
@@ 2.5 Computing the Radial Distribution Function
where $\psi(r)$ represents the summed number of particles located at a separation $r$ from each trial particle (the trial particles are taken to be the entire population of NN particles, double counting is allowed). The quantity $\delta V_{r}/V$ denotes the ratio of a small sampling volume at distance $r$ (i.e., a spherical shell of thickness $\delta r$) to the total measurement volume $V$.